OpenStudy (anonymous):

solve the system any algebraic method. 0.5x-0.3y=1.3 -1.4x+1.2y=-2.2

OpenStudy (anonymous):

(5, 4) Would you like an explanation?

OpenStudy (anonymous):

I think you should know how to do these by now, you've been giving them all day.

OpenStudy (anonymous):

make your life easy and start with \[5x-3y=13\] \[-14x+12y=-22\]

OpenStudy (anonymous):

multiply the first equation by 4 and add to second, get \[20x-12y=52\] \[-14x+12y=-22\] add to get \[6x=30\] so \[x=5\]

OpenStudy (anonymous):

then go back and replace x by 5 in either equation. .004*36i choose the first one \[5\times 5-3y=13\] \[25-3y=13\] \[-3y=-12\] \[y=4\]